OpenStudy (anonymous):

help someone plse: 2-a-4/a+3 = a^2-2/a+3

OpenStudy (anonymous):

(2-a-4)/(a+3) =(a^2-2)/(a+3) is this what you are asking jazi1?

OpenStudy (anonymous):

yes but nothing is in parenthese

OpenStudy (anonymous):

oh so it just looks like that on the question paper? because it could mean 2-a -(4/a+3) or [(2-a-4)/a] +3 and so on

OpenStudy (anonymous):

no it is 2 - a-4/a+3 = a^2-2/a+3 the 2 and - sign is in line with the bar that separates the left and right equation

OpenStudy (anonymous):

right then you would solve using the cross method

OpenStudy (anonymous):

coul you show me how plse

jhonyy9 (jhonyy9):

a-4 a^2 -2 2- ---- = ------- ------ 2(a+3) - (a-4) = a^2 -2 a+3 a+3 2a+6-a+4 = a^2 -2 a+10=a^2 -2 a^2 -a -12 =0 1+/- radical ( 1+48) 1+/- radical 49 1 +/- 7 a_1,_2 = ------------------- = ---------------- = -------- 2 2 2 a_1 = 8/2 =4 a_2 = -6/2 =-3 a_1 =4 a_2 =-3

OpenStudy (anonymous):

\[2-\frac{a-4}{a+3}=\frac{a^2-2}{a+3}\] Add the terms on the left side of the equals so as to get everything over (a+3) and then you can get rid of the (a+3)'s on both sides.